A DIVIDED IS TO BE SHARED AMONG THREE PEOPLE IN THE RATIO 1/2, 1/6, 1/3, IF THE THRID PERSON IS GIVEN $1100 HOW , MUCH DID THE FIRST TWO GET?
Answer by rfer(16322)

1100/(1/3)=
1100*3/1=3300
3300*1/2=1650
3300*1/6=550